# Post-detection inference for sequential changepoint localization

### v2

@@ 6 Conclusion
We developed a framework to construct a confidence set for the unknown changepoint $T$ after a sequential detection algorithm declares a change at a stopping time, providing conditional coverage guarantees (conditioned on the event that the stopping occurs after the changepoint) when the pre- and post-change distributions are known. We also extended our framework to handle composite pre- and post-change distributions and provide conditional coverage guarantees, demonstrating its good empirical performance. We impose no specific conditions on the change detection algorithms; they can function as entirely black-box or heuristic methods; we only need to be able to run the algorithm on simulated data sequences. By filling this significant gap in post-detection analysis, our work provides crucial tools for precise and timely changepoint localization in real-time data streams.
@@ Appendix A Omitted proofs
Here, we provide the formal mathematical details that were omitted in the main paper.
A.1 Omitted proofs from Section 2
Proof of Theorem 2.2.
First, observe that $M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}$ are i.i.d. under $H_{0,t}$.
[…]
$\mathbb{P}_{t}(M_{t}\leq\operatorname{Quantile}(1-\alpha;M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}))\geq\frac{\lceil(1-\alpha)(B+1)\rceil}{B+1}\geq 1-\alpha.$ (A.1)
For $L\in\mathbb{N}\cup\{\infty\},t\in\mathbb{N}$, define $C_{t,L}=\operatorname{Quantile}(1-\alpha;M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B}))$. Note that
$\mathbb{P}_{T}(T\in\{t\geq 1:M_{t}\leq C_{t,L}\},\tau<T)\leq\mathbb{P}_{T}[\tau<T]=\mathbb{\mathbb{P}}_{\infty}(\tau<T),$
the last equality holds because the event $[\tau<T]$ does not depend $X_{i}$ for $i\geq T$.
Also, observe that for any $L\in\mathbb{N}\cup\{\infty\}$, it follows from the definitions that $M_{t,L}^{j}\geq M_{t,\infty}^{j}$ and hence, $C_{t,L}\geq C_{t,\infty}$, for all $t\in\mathbb{N}$.
$1-\alpha$ $\leq\mathbb{P}_{T}(M_{T}\leq C_{T,\infty})$
$\leq\mathbb{P}_{T}(M_{T}\leq C_{t,L})$
$=\mathbb{P}_{T}(T\in\{t\geq 1:M_{t}\leq C_{t,L}\})$
$=\mathbb{P}_{T}(T\in\{t\geq 1:M_{t}\leq C_{t,L}\},\tau\geq T)+\mathbb{P}(T\in\{t\geq 1:M_{t}\leq C_{t,L}\},\tau<T)$
$\leq\mathbb{P}_{T}(T\in\mathcal{C})+\mathbb{\mathbb{P}}_{\infty}(\tau<T).$
Therefore, $\mathbb{P}_{T}(T\in\mathcal{C})\geq 1-\alpha-\mathbb{\mathbb{P}}_{\infty}(\tau<T)$.
Now, if $\mathcal{A}$ controls PFA at level $\delta$, i.e., $\mathbb{P}_{\infty}(\tau<\infty)\leq\delta$, we have
$\mathbb{P}_{T}(T\in\mathcal{C})\geq 1-\alpha-\mathbb{\mathbb{P}}_{\infty}(\tau<T)\geq 1-\alpha-\mathbb{\mathbb{P}}_{\infty}(\tau<\infty)\geq 1-\alpha-\delta.$
Proof of Proposition 2.4.
@@ A.1 Omitted proofs from Section 2
Note that $\mathcal{C}^{*}=\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq C_{t,L}\}\cup\{\tau+1,\tau+2,\cdots\}=\{t\geq 1:M_{t}\leq C_{t,L}\}$, since we defined $M_{t}=-\infty\leq C_{t,L}$, for $t>\tau.$
@@
$\mathbb{P}_{T}(T\in\mathcal{C}^{*})=\mathbb{P}_{T}(T\in\{t\geq 1:M_{t}\leq C_{t,L}\})=\mathbb{P}_{T}(M_{T}\leq C_{t,L})\geq\mathbb{P}_{T}(M_{T}\leq C_{T,\infty})\geq 1-\alpha,$
where the last inequality follows from (A.1). ∎
Proof of Corollary 2.6.
From Theorem 2.2, we get $\mathbb{P}_{T}(T\in\mathcal{C})\geq\mathbb{P}_{\infty}(\tau\geq T)-\alpha.$
Also, $\mathbb{P}_{T}(\tau\geq T)=\mathbb{P}_{\infty}(\tau\geq T)$ because of the fact the event $[\tau\geq T]$ does not depend $X_{i}$ for $i\geq T$. Therefore,
$\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)=\frac{\mathbb{P}_{T}(T\in\mathcal{C})}{\mathbb{P}_{T}(\tau\geq T)}\geq 1-\frac{\alpha}{\mathbb{P}_{\infty}(\tau\geq T)}.$
Therefore, if $\mathcal{A}$ controls PFA at level $\delta$, we have
$\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\frac{\alpha}{\mathbb{P}_{\infty}(\tau\geq T)}\geq 1-\frac{\alpha}{\mathbb{P}_{\infty}(\tau=\infty)}\geq 1-\frac{\alpha}{1-\gamma}.$
A.2 Omitted proofs from Section 3
@@
  1. (a)
    For $\mathcal{C}_{1}$: For all $t$, with $r_{t}=\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$, since $r_{t}$ is independent to $\{X_{n}\}_{n}$, we have
@@
  2. (b)
    For $\mathcal{C}_{2}$: Here, we have $r_{t}=1$. The proof for $\mathcal{C}_{2}$ is similar. For fixed $\alpha\in(0,1),$ with $\mathcal{S}^{\prime}_{\tau-t+1}$ having coverage $1-\beta$ we show that
    $P_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\alpha})\leq\alpha+\beta.$
    Then, we have
    $P_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{T,L}^{\alpha})\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+1-\mathbb{P}_{\theta_{0},\infty}(\tau\geq T).$
    If $\mathcal{A}$ controls PFA at level $\delta\in(0,1)$, we obtain
    $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\geq 1-\frac{\alpha+\beta}{\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)}\geq 1-\frac{\alpha+\beta}{\mathbb{P}_{\theta_{0},\infty}(\tau=\infty)}\geq 1-\frac{\alpha+\beta}{1-\delta}.$
    And similarly,
    $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2})\geq\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)-(\alpha+\beta)\geq 1-\alpha-\beta-\delta.$
A.3 Omitted proofs from Section 4
Proof of Theorem 4.2.
@@ A.3 Omitted proofs from Section 4
  1. (a)
    For $\mathcal{C}_{1}$: Also, for all $t$, with $r_{t}=\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$, since $r_{t}$ is independent to $\{X_{n}\}_{n}$, we have
    $\geq\mathbb{E}(1-\beta r_{t})=1-\beta\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t).$
@@
    $\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{o}\in\mathcal{S}_{t-1})$ $=\mathbb{E}(\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma r_{t})\mid r_{t}))$
    $\geq\mathbb{E}(1-\gamma r_{t})=1-\gamma\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t).$
@@
    $\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\alpha r_{t}})$
    $\leq\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\alpha r_{t}},\theta_{0}\in\mathcal{S}_{t-1},\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1})+\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{0}\notin\mathcal{S}_{t-1}))+\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\notin\mathcal{S}^{\prime}_{\tau-t+1})$
    $\leq\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>\operatorname{Quantile}(1-\alpha r_{t};M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B}))+(\beta+\gamma)\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$
    $\leq\mathbb{E}\left(\mathbb{P}_{\theta_{0},t,\theta_{1}}\left(M_{t}>\operatorname{Quantile}(1-\alpha r_{t};M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B})\mid r_{t}\right)\right)+(\beta+\gamma)\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
    $\leq\mathbb{E}(\alpha r_{t})+(\beta+\gamma)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$

### v3

@@ 7 Conclusion
We developed a general framework for constructing confidence sets for the unknown changepoint $T$ after a sequential detector signals a change at a data-dependent stopping time. It applies to any general nonparametric post-change class when the pre-change distribution is known, and we further extend it to settings with a composite pre-change class under a suitable assumption. For the parametric setting, we proposed an alternative simulation-based method that empirically yields tighter confidence sets when both pre- and post-change distributions are known and can handle dependent data as well. However, we note that our primary (universal) method is nonparametric, computationally more efficient, and simpler to implement. Crucially, all our methods are agnostic to the detection algorithm used and provide non-asymptotic (conditional) coverage guarantees, whereas prior works are tailored for exponential family distributions with the CUSUM procedure and provide only asymptotic (conditional) coverage. By bridging a longstanding gap in the literature, our work offers the first nonparametric, theoretically principled, and broadly applicable method for post-detection inference.
Supplementary Material
Outline of Supplementary Material
Omitted proofs can be found in Appendix A. We provide the implementation details of Algorithms 2 and 3 in Appendices B and C respectively, along with two concrete examples in both sections. Experimental results for the confidence intervals for pre- and post-change parameters are in Appendix D. Experimental results for a different implementation method are in Appendix E. Simpler algorithms for PFA-controlling change detectors are presented in Appendix F. Some additional experiments with dependent (Markov) pre- and post-change data, with detection procedures with bounded PFA, and with Poisson rate change problems are in Appendix G. Some heuristic algorithms are presented in Appendices H and I, pointing at open avenues for further theoretical work.
@@ Appendix A Omitted proofs
In this section, we provide the mathematical details that were omitted in the main paper.
Proof of Proposition 2.1..
@@
First note the fact that $\mathbb{P}_{F_{0},T,F_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau<T),$ since the event $[\tau<T]=\displaystyle\cup_{i=1}^{T-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq T\}$.
Let $\mathcal{C}\subseteq\{1,\dots,\tau\}$ be a confidence set satisfying $\mathbb{P}_{F_{0},t,F_{1}}(t\in\mathcal{C}\mid\tau\geq t)\geq 1-\alpha$, for all $t\in\mathbb{N}$. Since $\tau$ be a stopping time, $[\tau<t]=\displaystyle\cup_{i=1}^{t-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq t\}$. This implies $\mathbb{P}_{F_{0},\infty}(\tau\geq t)=\mathbb{P}_{F_{0},\infty}(\tau\geq t).$ Therefore,
$\mathbb{P}_{F_{0},t,F_{1}}(\phi^{(t)}(X_{1},\cdots,X_{\tau})=1)$ $=\mathbb{P}_{F_{0},t,F_{1}}(t\notin\mathcal{C})$
$=1-\mathbb{P}_{F_{0},t,F_{1}}(t\in\mathcal{C})$
$=1-\mathbb{P}_{F_{0},t,F_{1}}(t\in\mathcal{C}\mid\tau\geq t)\mathbb{P}_{F_{0},t,F_{1}}(\tau\geq t)$
$\leq 1-(1-\alpha)\mathbb{P}_{F_{0},\infty}(\tau\geq t)$
$\leq\alpha\mathbb{P}_{F_{0},\infty}(\tau\geq t),\leavevmode\nobreak\ \forall F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}.$
For the converse part, $\phi^{(t)}$ satisfies $\mathbb{P}_{F_{0},t,F_{1}}(\phi^{(t)}(X_{1},\cdots,X_{\tau})=1)\leq\alpha\cdot\mathbb{P}_{F_{0},\infty}(\tau\geq t),\forall F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}$
So, $\mathcal{C}=\{t\in\mathbb{N},t\leq\tau:\phi^{(t)}(X_{1},\cdots,X_{\tau})=0\}$ satisfies
$\mathbb{P}_{F_{0},T,F_{1}}(\phi^{(T)}(X_{1},\cdots,X_{\tau})=0)$
$=\mathbb{P}_{F_{0},T,F_{1}}(\phi^{(T)}(X_{1},\cdots,X_{\tau})=0,\tau\geq T)+\mathbb{P}_{F_{0},T,F_{1}}(\phi^{(T)}(X_{1},\cdots,X_{\tau})=0,\tau<T)$
$\leq\mathbb{P}_{F_{0},T,F_{1}}(T\in\{t\in\mathbb{N},t\leq\tau:\phi^{(t)}(X_{1},\cdots,X_{\tau})=0\})+\mathbb{P}_{F_{0},\infty}(\tau<T)$
$=\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})+\mathbb{P}_{F_{0},\infty}(\tau<T)$
Therefore, $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})\geq 1-\alpha\cdot\mathbb{P}_{F_{0},\infty}(\tau\geq T)-\mathbb{P}_{F_{0},\infty}(\tau<T)$, which implies
$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\}\mid\tau\geq T)=\frac{\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C},\tau\geq T)}{\mathbb{P}_{F_{0},T,F_{1}}(\tau\geq T)}=\frac{\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})}{\mathbb{P}_{F_{0},\infty}(\tau\geq T)}\geq 1-\alpha.$
Lemma A.1 (Ville’s inequality [32]).
If $\{R_{n}\}_{n\in\mathbb{N}}$ is an e-process for $\mathcal{P}$ with respect to a filtration $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$, then
$\mathbb{P}\left(\exists n\in\mathbb{N}:R_{n}\geq 1/\alpha\right)\leq\alpha,\forall\mathbb{P}\in\mathcal{P}.$ (A.1)
For a short proof, see [11][Lemma 1].
Proof of Theorem 3.1.
Note that by definition of $M_{t}$, we have $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq 2/\alpha,t=\hat{T}\leq\tau<\infty]=0$, $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq 2/\alpha,t>\tau]=0$ and $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}>2/\alpha,\tau=\infty]=0$. Therefore, the only cases where $M_{t}>2/\alpha$ might occur are $t<\hat{T}\leq\tau<\infty$ or $\hat{T}<t\leq\tau<\infty$. Hence, for any $t\in\mathbb{N}$,
$\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq 2/\alpha]$
$=\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq 2/\alpha,t<\hat{T}\leq\tau<\infty]+\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq 2/\alpha,\hat{T}<t\leq\tau<\infty]$
$=\mathbb{P}_{F_{0},t,F_{1}}\left[R^{(t)}_{\hat{T}-1}\geq 2/\alpha,t<\hat{T}\leq\tau<\infty\right]+\mathbb{P}_{F_{0},t,F_{1}}\left[S^{(t)}_{\hat{T}}\geq 2/\alpha,\hat{T}<t\leq\tau<\infty\right]$
$\leq\mathbb{P}_{F_{0},t,F_{1}}\left[R^{(t)}_{t\vee(\hat{T}-1)}\geq 2/\alpha\right]+\mathbb{P}_{F_{0},t,F_{1}}\left[S^{(t)}_{(t-1)\wedge\hat{T}}\geq 2/\alpha\right]$
Now, from Ville’s inequality (Lemma A.1) on the e-processes $\{R^{(t)}_{n}\}_{n\geq t}$,
$P_{F_{1}}\left(\exists n\geq t,R^{(t)}_{n}\geq 2/\alpha\right)\leq\alpha/2.$
Therefore,
$\mathbb{P}_{F_{0},t,F_{1}}\left[R^{(t)}_{t\vee(\hat{T}-1)}\geq 2/\alpha\right]$ $=\mathbb{P}_{F_{0},t,F_{1}}\left(\cup_{n=t}^{\infty}\left[R^{(t)}_{t\vee(\hat{T}-1)}\geq 2/\alpha,t\vee(\hat{T}-1)=n\right]\right)$
$\leq\mathbb{P}_{F_{0},t,F_{1}}\left(\cup_{n=t}^{\infty}\left[R^{(t)}_{n}\geq 2/\alpha\right]\right)$
$=P_{F_{1}}\left(\exists n\geq t,R^{(t)}_{n}\geq 2/\alpha\right)\leq\alpha/2.$
Similarly, for the e-processes $\{S^{(t)}_{n}\}_{n<t}$, we get $\mathbb{P}_{F_{0},t,F_{1}}\left[S^{(t)}_{(t-1)\wedge\hat{T}}\geq 2/\alpha\right]\leq\alpha/2.$ Therefore, from the above steps, we finally have
$\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq 2/\alpha]\leq\alpha/2+\alpha/2=\alpha.$
Proof of Theorem 3.2.
From Theorem 3.1, we have $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}}\mid r_{t}]\leq\alpha r_{t}$, since $r_{t}$ and $M_{t}$ are independent of each other. Hence,
$\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}}]=\mathbb{E}[\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}}\mid r_{t}]]\leq\mathbb{E}[\alpha r_{t}]=\alpha\mathbb{P}_{F_{0},\infty}[\tau\geq t],\forall t\in\mathbb{N}.$ Now,
$\mathbb{P}_{F_{0},T,F_{1}}(M_{T}<2/(\alpha r_{T}))$
$=\mathbb{P}_{F_{0},T,F_{1}}(T\in\{t\in\mathbb{N}:M_{t}<2/(\alpha r_{t})\})$
$\leq\mathbb{P}_{F_{0},T,F_{1}}(T\in\{t\in\mathbb{N}:M_{t}<2/(\alpha r_{t})\},\tau\geq T)+\mathbb{\mathbb{P}}_{F_{0},T,F_{1}}(\tau<T)$
$=\mathbb{P}_{F_{0},T,F_{1}}(T\in\{t\in\mathbb{N}:t\leq\tau,M_{t}<2/(\alpha r_{t})\}\mid\tau\geq T)\mathbb{P}_{F_{0},T,F_{1}}(\tau\geq T)+\mathbb{\mathbb{P}}_{F_{0},T,F_{1}}(\tau<T)$
$=\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)+1-\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T),$
where we twice used the fact that $\mathbb{P}_{F_{0},T,F_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau<T)$, which holds because $\tau$ is a stopping time and so the event $[\tau<T]=\displaystyle\cup_{i=1}^{T-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq T\}$. So, we obtain
$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\frac{\mathbb{P}_{F_{0},T,F_{1}}(M_{T}>2/(\alpha r_{T}))}{\mathbb{P}_{F_{0},\infty}(\tau\geq T)}\geq 1-\alpha.$ (A.2)
Proof of Proposition 3.4.
We only prove the case when $\theta_{0}^{*}=\sup\Theta_{0}\leq\inf\Theta_{1}$ and $\tau$ is the CUSUM type stopping time $\tau_{\text{wcs-ripr}}$ defined in (C.12). Proofs for the other cases, when $\theta_{0}^{*}=\inf\Theta_{0}\geq\sup\Theta_{1}$ and/or $\tau$ is the SR-type stopping time (C.13), follow analogously.
$\mathbb{P}_{\theta,\infty}[\tau_{\text{wcs-ripr}}\geq t]$ $=\mathbb{P}_{\theta,\infty}\left[\displaystyle\max_{1\leq k\leq t}\max_{1\leq j\leq k}\int_{\theta_{1}\in\Theta_{1}}\prod_{i=j}^{k}\frac{f_{\theta_{1}}(X_{i})}{f_{\theta_{0}^{*}}(X_{i})}dW(\theta_{1})<A\right].$
Since the family $\{F_{\theta}\}$ satisfies the MLR property in the statistic $T(x)$, and since $\frac{f_{\theta_{1}}(x)}{f_{\theta_{0}^{*}}(x)}$ is non-decreasing in $T(x)$, it follows that the random variables $\frac{f_{\theta_{1}}(X)}{f_{\theta_{0}^{*}}(X)}$ are stochastically increasing in $\theta$. That is, for $X\sim F_{\theta}$ and $X^{\prime}\sim F_{\theta^{\prime}}$ with $\theta\leq\theta^{\prime}\leq\sup\Theta_{0}$, we have $T(X)\preceq_{\text{st}}T(X^{\prime})$ (see e.g., [16, Chapter 3.4]) and hence,
$\frac{f_{\theta_{1}}(X)}{f_{\theta_{0}^{*}}(X)}\preceq_{\text{st}}\frac{f_{\theta_{1}}(X^{\prime})}{f_{\theta_{0}^{*}}(X^{\prime})}.$
By Theorem 1.A.2 of [24], the stochastic ordering implies that there exist a random variable $Z_{\theta_{1}}$ and functions $\psi_{\theta_{1},1}$ and $\psi_{\theta_{1},2}$ such that $\psi_{\theta_{1},1}(z)\leq\psi_{\theta_{1},2}(z)$ for all $z$ and $\frac{f_{\theta_{1}}(X)}{f_{\theta_{0}^{*}}(X)}\stackrel{{\scriptstyle d}}{{=}}\psi_{\theta_{1},1}(Z_{\theta_{1}})$ and $\frac{f_{\theta_{1}}(X^{\prime})}{f_{\theta_{0}^{*}}(X^{\prime})}\stackrel{{\scriptstyle d}}{{=}}\psi_{\theta_{1},2}(Z_{\theta_{1}})$. Therefore,
$\mathbb{P}_{\theta,\infty}[\tau_{\text{wcs-ripr}}\geq t]$ $=\mathbb{P}\left[\displaystyle\max_{1\leq k\leq t}\max_{1\leq j\leq k}\int_{\theta_{1}\in\Theta_{1}}\prod_{i=j}^{k}\psi_{\theta_{1},1}(Z_{\theta_{1},i})dW(\theta_{1})<A\right]$
$\geq\mathbb{P}\left[\displaystyle\max_{1\leq k\leq t}\max_{1\leq j\leq k}\int_{\theta_{1}\in\Theta_{1}}\prod_{i=j}^{k}\psi_{\theta_{1},2}(Z_{\theta_{1},i})dW(\theta_{1})<A\right]$
$=\mathbb{P}_{\theta^{\prime},\infty}[\tau_{\text{wcs-ripr}}\geq t].$
Proof of Theorem 3.5.
For all $t\in\mathbb{N},\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}^{*}}]=\mathbb{E}[\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}^{*}}\mid r_{t}]]\leq\mathbb{E}[\alpha r_{t}^{*}]=\alpha\mathbb{P}_{F_{0}^{*},\infty}[\tau\geq t]\geq\alpha\mathbb{P}_{F_{0},\infty}[\tau\geq t].$ The last step follows from Assumption 3.3. The rest of the proof is the same as in the proof of Theorem 3.2. ∎
Proof of Theorem 4.1.
First note the fact that $\mathbb{P}_{F_{0},T,F_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau<T),$ since the event $[\tau<T]=\displaystyle\cup_{i=1}^{T-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq T\}$.
Defining $C_{t,L}=\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})$, and $\hat{\alpha}_{t}=\alpha r_{t}$. Now,
$\mathbb{P}_{F_{0},T,F_{1}}(M_{T}\leq C_{T,L})$
$=\mathbb{P}_{F_{0},T,F_{1}}(T\in\{t\in\mathbb{N}:M_{t}\leq C_{t,L}\})$
$=\mathbb{P}_{F_{0},T,F_{1}}(T\in\{t\in\mathbb{N}:M_{t}\leq C_{t,L}\},\tau\geq T)+\mathbb{P}_{F_{0},T,F_{1}}(T\in\{t\in\mathbb{N}:M_{t}\leq C_{t,L}\},\tau<T)$
$\leq\mathbb{P}_{F_{0},T,F_{1}}(T\in\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq C_{t,L}\}\mid\tau\geq T)\mathbb{P}_{F_{0},T,F_{1}}(\tau\geq T)+\mathbb{\mathbb{P}}_{F_{0},T,F_{1}}(\tau<T)$
$=\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)+1-\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T),$
where we twice invoked $\mathbb{P}_{F_{0},T,F_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau<T)$ as explained earlier. So, we obtain
$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\frac{P_{T}(M_{T}>C_{T,L})}{P_{\infty}(\tau\geq T)}.$ (A.3)
First, observe that $M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}$ are i.i.d. under $H_{0,t}$. Hence, $M_{t}\leq\operatorname{Quantile}(1-c;M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B})$ $\iff M_{t}$ is one among the $\lceil(1-c)(B+1)\rceil$ smallest of $M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}$. So, for any fixed $c\in(0,1)$,
$\mathbb{P}_{F_{0},t,F_{1}}(M_{t}\leq\operatorname{Quantile}(1-c;M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}))\geq\frac{\lceil(1-c)(B+1)\rceil}{B+1}\geq 1-c\text{ for all }t.$
Since $\hat{\alpha}_{t}$ is independent of $M_{t}$, it follows that
$\mathbb{P}_{F_{0},t,F_{1}}(M_{t}\leq C_{t,\infty}\mid\hat{\alpha}_{t})\geq\frac{\lceil(1-\hat{\alpha}_{t})(B+1)\rceil}{B+1}\geq 1-\hat{\alpha}_{t},\text{ almost surely for all }t.$
@@
$\mathbb{P}_{F_{0},t,F_{1}}(M_{t}\leq C_{t,\infty})=\mathbb{E}_{t}(\mathbb{P}_{t}(M_{t}\leq C_{t,\infty}\mid\hat{\alpha}_{t}))\geq\mathbb{E}_{t}(1-\hat{\alpha}_{t})=1-\alpha\mathbb{P}_{F_{0},t,F_{1}}(\tau\geq t).$
As noted earlier, $\mathbb{P}_{F_{0},t,F_{1}}(\tau\geq t)=\mathbb{P}_{F_{0},\infty}(\tau\geq t)$ and hence, $\mathbb{P}_{F_{0},t,F_{1}}(M_{t}\leq C_{t,\infty})\geq 1-\alpha\mathbb{P}_{F_{0},\infty}(\tau\geq t).$ Also, for all $L\in\mathbb{N}$, it follows from the definitions that $M_{t,L}^{j}\geq M_{t,\infty}^{j}$. Therefore, $C_{t,L}\geq C_{t,\infty}$. Thus, for all $t$,
$\mathbb{P}_{F_{0},t,F_{1}}(M_{t}>C_{t,L})\leq\mathbb{P}_{F_{0},t,F_{1}}(M_{t}>C_{t,\infty})\leq\alpha\mathbb{P}_{F_{0},\infty}(\tau\geq t)$ (A.4)
Therefore, combining (A.3) and (A.4), we obtain $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha.$
@@
For all $t$, with $r_{t}=\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$, since $r_{t}$ is independent to $\{X_{n}\}_{n}$, we have
@@
Proof of Theorem 4.5..
@@
Also, for all $t$, with $r_{t}^{*}=\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$, since $r_{t}^{*}$ is independent to $\{X_{n}\}_{n}$,
$\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1})$ $=\mathbb{E}(\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t}^{*})\mid r_{t}^{*}))$
$\geq\mathbb{E}(1-\beta r_{t}^{*})=1-\beta\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t).$
@@
$\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{o}\in\mathcal{S}_{t-1})$ $=\mathbb{E}(\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma r_{t}^{*})\mid r_{t}^{*}))$
$\geq\mathbb{E}(1-\gamma r_{t}^{*})=1-\gamma\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t).$
@@
$\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\alpha r_{t}^{*}})$
$\leq\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\alpha r_{t}^{*}},\theta_{0}\in\mathcal{S}_{t-1},\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1})+\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{0}\notin\mathcal{S}_{t-1}))+\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\notin\mathcal{S}^{\prime}_{\tau-t+1})$
$\leq\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>\operatorname{Quantile}(1-\alpha r_{t}^{*};M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B}))+(\beta+\gamma)\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$
$\leq\mathbb{E}\left(\mathbb{P}_{\theta_{0},t,\theta_{1}}\left(M_{t}>\operatorname{Quantile}(1-\alpha r_{t}^{*};M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B})\mid r_{t}^{*}\right)\right)+(\beta+\gamma)\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
$\leq\mathbb{E}(\alpha r_{t}^{*})+(\beta+\gamma)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$